Project

General

Profile

NDOS Commissioning Meeting 2012/03/19

Attendance: Jon, Jan, Susan, Mark, Gavin, Kurt, Tomas, Minerba, Tom, Jarek, Peter

Agenda:

  1. Beam and Detector Update
  2. Cosmic Rate vs. Time in NDOS (Gavin)
  3. DataCheck Report (Jan)
  4. Installation Plans for the week
  5. DAQ Update and Plans for the week
  6. DCS Update and Plans for the week
  7. AOB

Beam and Detector Report

  • New DCMs and TDUs were installed last week, along with some new APDs. Several problems were encountered, but by the end of the week we got the full detector reading out with a mix of old and new hardware.
  • Peter: impressive that we got two completely incompatible systems working together in the same system.
  • Peter: the ~400 kHz noise on one DCM that exhibited noise was found to disappear
  • We cannot testing the timing synchronization across two master TDUs (we will never be able to). Peter is modifying the script to monitor the synch., but shifters need to be looking at the event display to make sure that the two TDUs are in synch.
  • Plan is to continue to take data with the two different master TDUs.

Understanding the Cosmic Rate Fluctuations (Gavin)

  • The ratio of Cosmic Rate to Active “good” channels seems to be a good handle on showing up anomalies
  • This should probably be added to the Nearline metrics for shifters to monitor deviations from the “norm”
  • Observing the event display is still the quickest way to catch things (you have to wait for the nearline metrics to be computed!)
  • Anomalies fall into a few categories:
    • Timing sync issues
    • DCMs (or perhaps channels in general) dropping in/out
    • Rate calculation requires track in both views (Biggest cause for deviations due to the detector configuration)
    • Not all time sync-isue runs removed from goodsynclist

DataCheck Report (Jan):

  • Didn't take much good data last week, so not much to report.
  • March 13, overnight run continuously lost channels.
  • Jon: pretty much all of last week was lost.
  • At end of week, a few runs had screwed up DCM mapping.
  • Nearline plots are currently empty since we did not collect data these past few days.

Installation plans for this week

  • No new hardware, perhaps some replacements.

DAQ Update and Plans for next week

  • Gennadiy would like to continue DCS tests (pushing logic from GUI down into IOC). Jon: let's aim for Wed. and Thurs.

DCS Update and Plans for next week

  • No report

AOB

  • None